OpenStudy (anonymous):

Can someone please help me I am stuck on this!! Are a rational expression and its simplified forms equivalent? Provide mathematical examples to support your opinions. You may use tables or graphs to organize and present your thoughts.

Directrix (directrix):

@Pologirl19 I need help on what you mean by rational expression. Are we talking 16/12 is equivalent to 4/3? Or, are we wondering if y =[(x-1)(x+2)] /(x+2) is equivalent to y = (x - 1) ?

OpenStudy (anonymous):

The first one

Directrix (directrix):

Are we talking 16/12 is equivalent to 4/3? ---------------- Yes, these are equivalent. I am unsure as to what the exact question is.

OpenStudy (anonymous):

Okay thank you